What amount of current should flow through a solenoid of 330 turns and diameter 4.5cm, for it to produce a magnetic field strength of 400mT


1
Expert's answer
2021-07-06T08:23:22-0400

We can find the amount of current should flow through a solenoid as follows:


"B=\mu_0nI,""I=\dfrac{B}{\mu_0n}=\dfrac{400\cdot10^{-3}\ T}{4\pi\cdot10^{-7}\ \dfrac{T}{A\cdot m}\cdot330\ \dfrac{turns}{m}}=965\ A."